Solve the following formula for n and state any restrictions: l=a+(n-1)d; first letter is an L.
Answer by edjones(8007)

L=a+d(n-1)
-d(n-1)=a-L Subtract d(n-1)+L from each side of the equation.
d(n-1)=L-a Multiply -1 by each side of the equation.
n-1=(L-a)/d Divide d by each side of the equation.
n=(L-a)/d + 1 Add 1 to each side of the equation.
.
Ed